OpenStudy (bossimbacon):

Joshua paid $28 for a pair of shoes that has been discounted 20%. What is the original price of the shoes before the discount? @shadowlegendx

OpenStudy (seascorpion1):

Let the original price =x therefore: x-20%x=28 or 0.8x=28

OpenStudy (bossimbacon):

but that would make the answer 0.04 @seascorpion1

OpenStudy (seascorpion1):

28/0.8 can't be 0.04

OpenStudy (bossimbacon):

oh i mean 70 :)

OpenStudy (seascorpion1):

That makes sense :)

OpenStudy (seascorpion1):

Sorry (should've switched my brain on) 70 doesn't make sense at all. Can you check your calculation and let me know what your answer is? You should be able to plug it back into the original question to check that it's correct.

OpenStudy (shadowlegendx):

If x is the original price...it doesnʻt equal 28

OpenStudy (seascorpion1):

The original price (lets call it x) minus 20% of the original price must equal the new price (i.e. 28). So x-(20%*x)=28. Or x-0.2x=28.